# On the Quest for Effectiveness in Human Oversight: Interdisciplinary Perspectives

### v1

@@ 4.3. Environmental circumstances
Under “environmental circumstances” we subsume factors that lie outside the technical system or the human overseer, yet they can influence the effectiveness of oversight. This discussion could also include aspects such as stakes of the situation (Langer and Landers, 2021), single vs. multitasking environments (Lyons and Guznov, 2018), or whether oversight is organized as a team task (Gao et al., 2014; Ulfert et al., 2023).
@@
Adequate job design. The job of a human overseer needs to be well-designed (Parker and Grote, 2020) meaning that the job is, e.g., motivating, satisfying, and not overly demanding. For example, the job characteristics model [^8] See also other models from psychology that inspire job design to make the work more motivating and satisfying (e.g., (Humphrey et al., 2007)), or less demanding (e.g. the job-demands-resources model (Demerouti et al., 2001)). from work psychology (Hackman and Oldham, 1976) highlights that certain job characteristics promote motivation and satisfaction with the job, namely skill variety , task identity , perceived task significance , autonomy , and feedback from the job. The design of the job is important for promoting self-control, epistemic access, and fitting intentions, for example via motivation (cf. Section 4.2). Also, ill-designed jobs can lead to counterproductive work behavior, i.e., behavior at work that is inconsistent with or even intentionally disrespecting job duties (e.g., not even trying to maintain a sufficient epistemic access, or intentionally overwriting outputs to feel “needed” (Strich et al., 2021)).
@@
Role conflicts. Role conflicts could arise if the overseer is at the same time a decision-maker who uses the AI-based system to support their decisions (King and King, 1990). Consider, for example, a hiring scenario where an individual juggles the roles of both human overseer, tasked with mitigating unfairness, and an HR manager, who is responsible for inviting only the best candidates to a job interview. A conflict arises when the only method to address unfairness involves inviting additional, apparently less suitable candidates, conflicting with the HR manager’s goal of selecting the best candidates only. This misalignment of duties may give rise to intentions that are unfitting for an overseer.
@@
Independent thinking. Processing a situation under the impression of a system’s output can affect an overseer’s judgment of such an output (Furnham and Boo, 2011). For example, seeing a recidivism risk score for a defendant at court in addition to the information about the defendant’s case may affect how people judge this case. Seeing a “high anchor” (i.e., the system’s prediction of a high risk for recidivism) will make it more likely that the judge will end up deciding that the defendant has a high risk for recidivism. Such anchors make overseers selectively aware of anchor-consistent information (Englich et al., 2006) which undermines epistemic access.
@@
In the area of AI-supported decision-making, there are several ideas of how to possibly prevent such effects. For instance, although leading to less acceptance of the system, it seems to partly counter such anchoring effects if people are forced to think about a given case themselves for some time before they make a final decision (Buçinca et al., 2021). Alternatively, overseers may only receive the AI-based output after they have made up their own mind about a specific case because then they would not initially be affected by the AI-based output and may be in a better position to judge whether the AI-based output is appropriate (Green, 2022; Zerilli et al., 2019).
@@
Accountability. Here we refer to accountability as “an obligation to explain and to justify [one’s] conduct” ((Bovens, 2007), p.447). If overseers feel accountable for their actions this may influence their attentiveness or may lead to them feeling more stressed (Hall et al., 2015). A feeling of accountability could thus promote efforts of overseers to bring themselves in a sufficient epistemic position. It can also be hypothesized that a felt obligation to explain oneself will be motivating to maintain high self-control in order to keep up with the expectations of the job, and that it will incentivize to uphold intentions that fit one’s duties. However, it could also lead to unintended consequences such as leading to unfit intentions: if overseers believe they would have no good explanation for overriding a system output, they may not do so in cases where they know that they have to file a report after overwriting system outputs.
@@
Time pressure. Significant time pressure may undermine various conditions for effectiveness (Rieger and Manzey, 2022). For example, if an overseer only has a few seconds to react to warnings by the system, this could hinder sufficient epistemic access. In extreme cases, time pressure could also undermine an overseers’ causal power. For instance, if the option to overwrite an output is only available for split seconds, the overseer may not be able to press the button.
@@ 5. Application to the AI Act Proposal
Finally, our analysis turns to effective human oversight as outlined in the final version of the AI Act (European Parliament, 2024) [^9] At the time of writing, the trilogue negotiations between the European Commission, the Parliament and the Council have been concluded and a finalized version has been approved by the European Parliament. However, the final version of the AI Act has not yet been approved by the Council of the EU.. According to Article 14(1) of the proposal, what the AI Act defines as high-risk AI systems shall be designed and developed in such a way that they can be effectively overseen during their use. As stated in Article 14(2), this oversight shall aim to prevent or minimize risks to health, safety or fundamental rights. Risk mitigation is therefore defined as a key objective of human oversight measures which should contribute to the overall trustworthiness of high-risk systems (Enqvist, 2023). Consequently, the effectiveness of human oversight, as outlined in Article 14(1), serves as a foundational principle that requires oversight measures to be capable of achieving their risk mitigation objective. This principle can be understood as a condition for a minimum level of effectiveness of the oversight measures implemented, comparable to a threshold that must be reached.
@@ 5.1. Conditions for effectiveness in our proposal and the AI Act
Article 14(3) obliges the provider to enable its deployer to perform effective oversight by identifying “appropriate measures”. [^10] It should be noted that the Commission proposal (European Commission, 2021) referred to the deployer as the “user”. The European Parliament (European Parliament, 2023) rightly proposed to use the term “deployer” instead to avoid any confusion with the end user of the AI system. This proposal found its way in the version approved by the Parliament (European Parliament, 2024). We therefore use the term “deployer”. Article 14(4) contains certain aspects of what the proposed AI Act envisages as indicators of effectiveness of these measures. These require to enable the individuals to whom human oversight is assigned to do the following, as appropriate and proportionate to the circumstances:
@@
(b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (‘automation bias’), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
@@
(quoted from Article 14(4) of the final version approved by the European Parliament(European Parliament, 2024))
These measures reflect our conditions of effectiveness. Requirements (d) and (e) underscore the EU’s recognition of the importance of human overseers being able to influence the AI system, either by modifying the system’s output or by stopping its operation. These requirements manifest the need for both causal power and self-control, ensuring that individuals can influence the decisions of the AI system, or decide not to use the system’s output or to modify it. Conditions (a) and (c) relate to the understanding of the system, as they require the provider to enable the overseer to understand the limitations of the system (though not necessarily all details of the system), to monitor it effectively, and to interpret its results correctly. This is consistent with the concept of epistemic access, which enables the overseer to make informed decisions. Awareness of automation bias, as described in (b), is intertwined with epistemic access and self-control, as described above in Section 4.2. Furthermore, the conditions for effectiveness must also be implemented by the deployer. In this regard, Article 26(2) (European Commission, 2021) emphasizes the need for the necessary competence, training, authority, and support in oversight functions, hence requiring epistemic access and causal power.  [^11] This requirement was mentioned only in the non-binding recital (48) of the EU Commission’s proposal (European Commission, 2021), but was ultimately stipulated more directly in Article 29 of the proposals by the EU Council (Council of the European Union, 2022), the Parliament (European Parliament, 2023) and Article 26 of the approved version (European Parliament, 2024). The criterion of fitting intentions of the agents involved has little direct equivalent in the legal text, which is not focused on subjective intentions but on actions. The parties’ underlying intentions and motivations are not directly relevant. One might consider, however, that the automation bias targeted in Article 14(4) (b) may interfere with fitting intentions of the human overseer (cf. Section 4.2). Furthermore, the significant penalties according to Article 99 (European Parliament, 2024) strongly incentivize norm-compliant behavior.
@@ 5.2. Further analysis of the EU legislation’s approach
It is striking that the measures appearing in Article 14(4) are of varying degrees of abstraction: While some measures contain specific requirements for the design of the AI-system, other measures are less specific. For example, the requirement of a “stop” button in (e) is highly concrete. In contrast, the requirement in (a) to create an understanding of the limitations of the system is vague and leaves a lot of room for interpretation. Overall, the requirements for the various players – providers of the AI system, deployers, the person entrusted with oversight – could have been made clearer in the structure of the AI Act. Thus, Article 14(4) appears to be a loose collection of items that were deemed useful in the legislative process without much structural consistency. The lack of regulatory clarity could hinder the practical applicability of Article 14 and thus the introduction of truly effective human oversight.
@@
Measures in accordance with (b), which are intended to maintain an awareness of the automation bias, seem out of place: it introduces a concrete psychological aspect that sticks out from the more technical and operational aspects outlined in the other provisions, especially considering that automation bias is just one of many relevant biases (e.g., hindsight bias, confidence bias (Hilbert, 2012; Roese and Vohs, 2012)), heuristics (e.g., availability heuristic, (Tversky and Kahneman, 1973)), and psychological phenomena (e.g., exhaustion, motivation, (Ryan and Deci, 2000; Sonnentag et al., 2023)) that would need to be accounted for.
@@
Not every measure in Article 14(4) will have to be fully implemented in every high-risk system. This already follows from the restriction that the measures should be implemented “as appropriate and proportionate to the circumstances”. As a result, the specific measures taken must be assessed on a case-by-case basis. Not only how these requirements apply, but also which requirements are applicable therefore depends heavily on the individual case and the design of the system. While this approach provides flexibility that accommodates the sometimes beneficial vagueness of the effectiveness of human oversight also mentioned in Section 3.1.5, it may also be associated with greater legal uncertainty. The legislation will be complemented by technical standards developed by European standardization organisations to at least partially remove this legal uncertainty and to facilitate the implementation of requirements for high-risk systems, including human oversight (Veale and Borgesius, 2021). However, this standardization process is still in its early stages and is limited to a few general guidelines (Laux, 2023). Moreover, the question arises as to what requirements these standards must meet to satisfy the effectiveness requirement of Article 14(1). While Article 14(4) offers some guidance, its vagueness and varying levels of abstraction are insufficient for determining true effectiveness in specific cases. Thereby, the standardization organisations risk developing technical standards that may not withstand legal scrutiny. So, the EU’s reliance on the AI Act’s effective implementation through practical technical standards poses a potential threat to the Regulation’s requirement for human oversight.
@@ 5.3. Utilizing our definition of effectiveness
Our objective-first approach to effectiveness can be used to integrate the piecemeal list of different requirements listed in the AI Act into a coherent system. This is advantageous for a legally sound application of the AI Act, as it helps to structure the requirements for and responsibilities of the different parties involved. One of the benefits of the breadth and structure of our conceptualization is that it simplifies the assessment process for regulators, making it easier to evaluate different AI systems against a structured set of principles that are framed by the four conditions. The three domains of facilitators and inhibitors can aid in giving further structure for certain use cases. This streamlining may also decrease the legal uncertainty in Article 14, fostering a more consistent and predictable environment for AI providers and deployers alike, while providing a practical definition of effectiveness for deciding courts. This could also help to uncover further shortcomings in Article 14(4): For instance, while understanding the system’s output (c) can enhance epistemic access, understanding of the system’s input can be equally crucial. Our conditions for effectiveness, when applied with expertise, make this need evident; whereas this need can be challenging to identify when relying solely on Article 14(4), which does not mention the system’s input at all.
@@
Consequently, the points listed in Table 1 can be understood as examples of how an assessment of the effectiveness of measures could be carried out. Defining effectiveness using our four conditions provides a practical framework that highlights the key aspects of effective oversight. Our approach is therefore well suited to assist standard-setting organizations and to clarify the rather broad requirements of Article 14(4). In addition, our proposed framework provides guidance to practitioners who do not wish to rely solely on technical standards. Finally, our approach can assist courts in the future by providing a framework for defining and assessing the effectiveness of human oversight.
6. Limitations
@@
As discussed above, we see advantages brought about by our conceptualization of effective human oversight. Nevertheless, we also want to point out three limitations as well as future points to address. First, our conceptualization is designed for broad applicability across various AI-based systems, and thus shares a common criticism with the AI Act, namely that of residing on a high level of abstraction and that of vagueness, as discussed in Section 3.1.5. Dealing with these issues, especially in complex scenarios, requires (a) involving individuals with extensive expertise in the task at hand and (b) fostering collaboration in multidisciplinary teams to assess the impact on human oversight practice. Second, when we came up with the facilitators and inhibitors spelled out in Section 4, we realized that our conceptualization is helpful for thinking through conditions that could promote or hinder effective oversight. However, for some of the facilitators (or inhibitors), it depends on the concrete context which of the four conditions they contribute to (or hinder). Moreover, for certain cases, the mapping from facilitators and inhibitors to conditions needs considerable empirical research, outside the scope of this conceptual paper. Third, we did not discuss the concept of human oversight itself in much detail, but focused only effectiveness of human oversight. So, while our paper is helpful in determining whether an instance of human oversight is effective, it is less helpful in assessing whether something is human oversight. However, our definition of effectiveness is broad enough to be applied to many roles along the AI pipeline, not only to overseers. Some more conceptual discussion of human oversight itself is for instance provided by (Laux, 2023).

### v2

@@ 4.3. Environmental circumstances
Under “environmental circumstances”, we subsume factors that lie outside the technical system or the oversight person, yet they can influence the effectiveness of oversight. This discussion could also include aspects such as stakes of the situation (Langer and Landers, 2021), single vs. multitasking environments (Lyons and Guznov, 2018), or whether oversight is organized as a team task (Gao et al., 2014; Ulfert et al., 2023).
@@
Adequate job design. The job of an oversight person needs to be well designed (Parker and Grote, 2020), which means that the job is, e.g., motivating, satisfying, and not overly demanding. For example, the job characteristics model from work psychology (Hackman and Oldham, 1976) highlights that certain job characteristics promote motivation and satisfaction with the job, namely skill variety, task identity, perceived task significance, autonomy, and feedback from the job. [^8] See also other models from psychology that inspire job design to make the work more motivating and satisfying (e.g., (Humphrey et al., 2007)), or less demanding (e.g. the job-demands-resources model (Demerouti et al., 2001)). The design of the job is important for promoting self-control, epistemic access, and fitting intentions, for example via motivation (cf. Section 4.2). Also, ill-designed jobs can lead to counterproductive work behavior, i.e., behavior at work that is inconsistent with or even intentionally disrespecting job duties (e.g., not even trying to maintain sufficient epistemic access, or intentionally overwriting outputs to feel “needed” (Strich et al., 2021)).
@@
Role conflicts. Role conflicts could arise if the oversight person is at the same time a decision-maker who uses the AI-based system to support their decisions (King and King, 1990). Consider, for example, a hiring scenario where an individual juggles both the role of an oversight person tasked with mitigating unfairness and the role of an HR manager who is responsible for inviting only the best candidates to a job interview. A conflict arises when the only method to address unfairness involves inviting additional, apparently less suitable candidates, conflicting with the HR manager’s goal of selecting the best candidates only. This misalignment of duties may give rise to intentions that are unfitting for an oversight person.
@@
Independent thinking. Processing a situation under the impression of a system’s output can affect an oversight person’s judgment of such an output (Furnham and Boo, 2011). For example, seeing a recidivism risk score for a defendant in court, in addition to the information about the defendant’s case, may affect how people judge this case. Seeing a “high anchor” (i.e., the system’s prediction of a high risk for recidivism) will make it more likely that the judge will decide that the defendant has a high risk for recidivism. Such anchors make oversight persons selectively aware of anchor-consistent information (Englich et al., 2006) which undermines epistemic access.
@@
In the area of AI-supported decision-making, there are several ideas of how to possibly prevent such effects. For example, although this leads to less acceptance of the system, it seems to partly counter such anchoring effects if people are forced to think about a given case themselves for some time before making a final decision (Buçinca et al., 2021). Alternatively, oversight persons may only receive the AI-based output after they have made up their own mind about a specific case because then they would not initially be affected by the AI-based output and may be in a better position to judge whether the AI-based output is appropriate (Green, 2022; Zerilli et al., 2019).
@@
Accountability. Here we refer to accountability as “an obligation to explain and to justify [one’s] conduct” (Bovens, 2007, p.447). If oversight persons feel accountable for their actions this may influence their attentiveness or may lead to them feeling more stressed (Hall et al., 2015). Thus, a feeling of accountability could promote efforts of oversight persons to get themselves in a sufficient epistemic position. It can also be hypothesized that a felt obligation to explain oneself will motivate to maintain high self-control in order to keep up with the expectations of the job and that it will incentivize one to uphold intentions that fit one’s duties. However, it could also lead to unintended consequences such as leading to unfit intentions: if oversight persons believe that they would have no good explanation for overriding a system output, they may not do so in cases where they know that they have to file a report after overwriting system outputs.
@@
Time pressure. Significant time pressure may undermine various conditions for effectiveness (Rieger and Manzey, 2022). For example, if an oversight person only has a few seconds to react to warnings by the system, this could hinder sufficient epistemic access. In extreme cases, time pressure could also undermine an oversight persons’ causal power. For instance, if the option to overwrite an output is only available for split seconds, the oversight person may not be able to press the button.
@@ 5. Application to the AI Act Proposal
Finally, our analysis turns to effective human oversight as outlined in the AI Act (European Parliament, 2024a) [^9] At the time of writing, the trilogue negotiations between the European Commission, the Parliament, and the Council have been concluded and a finalized version has been approved by the European Parliament. However, the final version of the AI Act has not yet been approved by the Council of the EU.. According to Article 14(1) of the proposal, what the AI Act defines as high-risk AI systems shall be designed and developed in such a way that they can be effectively overseen during their use. As stated in Article 14(2), this oversight shall aim to prevent or minimize risks to health, safety, or fundamental rights. Risk mitigation is therefore defined as a key objective of human oversight measures that should contribute to the overall trustworthiness of high-risk systems (Enqvist, 2023). Consequently, the effectiveness of human oversight, as outlined in Article 14(1), serves as a foundational principle that requires oversight measures to be capable of achieving their risk mitigation objective. This principle can be understood as a condition for a minimum level of effectiveness of the oversight measures implemented, comparable to a threshold that must be reached.
@@ 5.1. Conditions for effectiveness in our proposal and the AI Act
Article 14(3) obliges the provider to enable its deployer to perform effective oversight by identifying “appropriate measures”. [^10] It should be noted that the Commission proposal (European Commission, 2021) referred to the deployer as “user”. The European Parliament (European Parliament, 2023) rightly proposed to use the term “deployer” instead to avoid any confusion with the end user of the AI system. This proposal found its way into the version approved by the Parliament (European Parliament, 2024a). We therefore use the term “deployer”. Article 14(4) contains certain aspects of what the proposed AI Act envisages as indicators of the effectiveness of these measures. These require enabling the individuals to whom human oversight is assigned to do the following, as appropriate and proportionate to the circumstances:
@@
(b) to remain aware of the possible tendency of automatically relying or over-relying on the output produced by a high-risk AI system (automation bias), in particular for high-risk AI systems used to provide information or recommendations for decisions to be taken by natural persons;
@@
(quoted from Article 14(4) of the amendment to the version approved by the European Parliament (European Parliament, 2024a))
These measures reflect our conditions of effectiveness. Requirements (d) and (e) underscore the EU’s recognition of the importance of oversight persons being able to influence the AI system, either by modifying the system’s output or by stopping its operation. These requirements manifest the need for both causal power and self-control, ensuring that individuals can influence the decisions of the AI system, or decide not to use the system’s output or to modify it. Conditions (a) and (c) relate to the understanding of the system, as they require the provider to enable the oversight person to understand the limitations of the system (though not necessarily all details of the system), to monitor it effectively, and to interpret its results correctly. This is consistent with the condition of epistemic access, which enables the oversight person to make informed decisions. Awareness of automation bias, as described in (b), is intertwined with epistemic access and self-control, as described above in Section 4.2. Furthermore, Article 26(2) (European Parliament, 2024a) emphasizes the need for the necessary competence, training, authority, and support in oversight functions, hence requiring epistemic access and causal power. [^11] This requirement was mentioned only in the non-binding recital (48) of the EU Commission’s proposal (European Commission, 2021), but was ultimately stipulated more directly in Article 29 of the proposals by the EU Council (Council of the European Union, 2022), the Parliament (European Parliament, 2023) and Article 26 of the approved version (European Parliament, 2024b). The criterion of fitting intentions of the agents involved has little direct equivalent in the legal text, which is not focused on subjective intentions but on actions. The parties’ underlying intentions and motivations are not directly relevant. However, one might consider that the automation bias targeted in Article 14(4) (b) may interfere with the fittingness of the oversight person’s intentions (cf. Section 4.2). Furthermore, the significant penalties according to Article 99 (European Parliament, 2024a) strongly incentivize norm-compliant behavior.
@@ 5.2. Further analysis of the EU legislation’s approach
It is striking that the measures appearing in Article 14(4) are of varying degrees of abstraction: While some measures contain specific requirements for the design of the AI system, other measures are less specific. For example, the requirement of a “stop” button in (e) is highly concrete. In contrast, the requirement in (a) to create an understanding of the limitations of the system is vague and leaves a lot of room for interpretation. Overall, the requirements for the various parties – providers of the AI system, deployers, and the person entrusted with oversight – could have been made clearer in the structure of the AI Act. Thus, Article 14(4) appears to be a loose collection of items that were deemed useful in the legislative process without much structural consistency. The lack of regulatory clarity could hinder the practical applicability of Article 14 and, thus, the introduction of truly effective human oversight.
@@
Measures in accordance with (b), which are intended to maintain an awareness of the automation bias, seem out of place: it introduces a concrete psychological aspect that sticks out from the more technical and operational aspects outlined in the other provisions, especially considering that automation bias is just one of many relevant biases (e.g., hindsight bias, confidence bias (Hilbert, 2012; Roese and Vohs, 2012)), heuristics (e.g., availability heuristic (Tversky and Kahneman, 1973)), and other psychological phenomena (e.g., exhaustion, motivation (Ryan and Deci, 2000; Sonnentag et al., 2023)) that would need to be accounted for.
@@
Not every measure in Article 14(4) will have to be fully implemented in every high-risk system. This already follows from the restriction that the measures should be implemented “as appropriate and proportionate to the circumstances”. As a result, the specific measures taken must be assessed on a case-by-case basis. Not only how these requirements apply, but also which requirements are applicable depends heavily on the individual case and the design of the system. Although AI Act’s approach provides flexibility that accommodates the sometimes beneficial vagueness of the effectiveness of human oversight also mentioned in Section 3.1.5, it may also be associated with greater legal uncertainty. The legislation will eventually be complemented by technical standards developed by European standardization organizations to at least partially remove this legal uncertainty and to facilitate the implementation of requirements for high-risk systems, including human oversight (Veale and Borgesius, 2021). However, this standardization process is still in its early stages and is limited to a few general guidelines (Laux, 2023). Moreover, the question arises as to what requirements these standards must meet to satisfy the effectiveness requirement of Article 14(1). While Article 14(4) offers some guidance, its vagueness and varying levels of abstraction are insufficient for determining true effectiveness in specific cases. Thereby, the standardization organisations risk developing technical standards that may not withstand legal scrutiny. So, the EU’s reliance on the AI Act’s effective implementation through practical technical standards poses a potential threat to the Regulation’s requirement for human oversight.
@@ 5.3. Utilizing our definition of effectiveness
Our objective-first approach to effectiveness can be used to integrate the piecemeal list of different requirements listed in the AI Act into a coherent system. This is advantageous for a legally sound application of the AI Act, as it helps to structure the requirements for and responsibilities of the different parties involved. One of the benefits of the breadth and structure of our conceptualization is that it simplifies the assessment process for regulators, making it easier to evaluate different AI systems against a structured set of principles that are framed by the four conditions. The three domains of facilitators and inhibitors can aid in giving further structure for certain use cases. This streamlining may also decrease the legal uncertainty in Article 14, fostering a more consistent and predictable environment for AI providers and deployers alike, while providing a practical definition of effectiveness for deciding courts. This could also help uncover further shortcomings in Article 14(4): For instance, while understanding the system’s output (c) can enhance epistemic access, understanding of the system’s input can be equally crucial. Our conditions for effectiveness, when applied with expertise, make this need evident; whereas this need can be challenging to identify when relying solely on Article 14(4), which does not mention the system’s input at all.
@@
Consequently, the points listed in Table 1 can be understood as examples of how an assessment of the effectiveness of measures could be carried out. Defining effectiveness using our four conditions provides a practical framework that highlights the key aspects of effective oversight. Our approach is, therefore, well suited to assist standard-setting organizations and to clarify the rather broad requirements of Article 14(4). In addition, our proposed framework provides guidance to practitioners who do not wish to rely solely on technical standards. Finally, our approach can assist courts in the future by providing a framework for defining and assessing the effectiveness of human oversight.
6. Limitations and Future Work
@@
As discussed above, we see advantages brought about by our conceptualization of effective human oversight. Nevertheless, we also want to point out three limitations as well as future points to address. First, our conceptualization is designed for broad applicability across various AI-based systems, and thus shares a common criticism with the AI Act, namely that of residing on a high level of abstraction and that of vagueness, as discussed in Section 3.1.5. Dealing with these issues, especially in complex scenarios, requires (a) involving individuals with extensive expertise in the task at hand and (b) fostering collaboration in multidisciplinary teams to assess the impact on human oversight practice. Second, when we came up with the facilitators and inhibitors spelled out in Section 4, we realized that our conceptualization is helpful for thinking through conditions that could promote or hinder effective oversight. However, for some of the facilitators (or inhibitors), the question of which of the four conditions they contribute to (or hinder) depends on the concrete context. Moreover, for certain cases, the mapping from facilitators and inhibitors to conditions needs considerable empirical research, outside the scope of this conceptual paper. Third, we did not discuss the concept of human oversight itself in much detail, but focused only effectiveness of human oversight. So, while our paper is helpful in determining whether an instance of human oversight is effective, it is less helpful in assessing whether something is human oversight. However, our definition of effectiveness is broad enough to be applied to many roles along the AI pipeline, not only to oversight persons. Some more conceptual discussion of human oversight itself is, for instance, provided by (Laux, 2023).
Future efforts will aim to enhance the practical applicability of our framework. We plan to develop a comprehensive practical framework, help to establish best practices, and create tools for implementing and auditing effective human oversight. These developments can benefit both policymakers and practitioners. This endeavor will require ongoing interdisciplinary collaboration, such as conducting empirical studies to validate and refine oversight mechanisms and designing holistic requirements for high-risk systems, oversight personnel, and their environments. Additionally, efforts should include creating educational programs and training for oversight persons, as well as engaging with regulatory bodies to ensure the ethical deployment of AI systems.